OpenStudy (anonymous):

What is the solution of the following system? -3x - -2y= -12 9x + 6y = -9

OpenStudy (anonymous):

is there a typo in this problem because i see 2 minus's? well anyways what you want to do is on the first equation solve for y then plug it into the 2nd equation and solve for x then plug that x value into the first equation and solve for y to get your answers

OpenStudy (anonymous):

one is a minus and one is a negative

OpenStudy (anonymous):

ok so when you subtract a negative it becomes a positive so your first equation is -3x + 2y = -12 9x + 6y = -9 multiply first eq by 3 -9x + 6y = -12 9x + 6y = -9 add them 12y = -21 y = -21/12 = -7/4 substitute in 1st equation -3x + 2(-7/4) = -12 -3x + -7/2 = -12 -3x = 17/2 x = -17/6

OpenStudy (anonymous):

(2, 1) no solutions (–2, –1) infinitely many solutions so than witch answer could it be?

OpenStudy (anonymous):

sorry i made a typo in my work -3x + 2y = -12 9x + 6y = -9 y = -6 + 3/2(x) 9x+6(-6+3/2(x)) = -9 9x - 36 + 9x = -9 18x = 27 x = 1.5 -3(1.5) + 2y = -12 y = -3.75 might of made a mistake neeed someone to double check